OpenStudy (anonymous):

Identify the vertex for the graph of y = -3x2 - 6x + 4.

OpenStudy (anonymous):

this is for a study guide and it is a multiple chocie question, the answer choices are (1,13) (-1,13) (1,7) or (-1,7)

OpenStudy (lalaly):

my bad one min

OpenStudy (anonymous):

okay :)

OpenStudy (shayaan_mustafa):

Hi gatorgalvet :) make right side of the question as a perfect square.

OpenStudy (lalaly):

\[y=-3[x^2+2x-\frac{4}{3}]\]\[y=-3[(x+1)^2-1-\frac{4}{3}]\]\[y=-3[(x+1)^2-\frac{7}{3}]\]\[y=-3(x+1)^2+7\] so vertex is (-1,7) i messed up witht he negatives

OpenStudy (zepp):

\(\large \frac{-b}{2a}\)!

OpenStudy (anonymous):

okay thx zepp, shay, and lalaly!

OpenStudy (zepp):

\(\huge \frac{-(-6)}{2(-3)}=\frac{6}{-6}=-1\) That's x, then plug it in the equation to find y, and you'll find what @lalaly got :]